suppose A and B are 3x3 matrices.Show that if B is obtained from A by adding 2 times the first row of A to the last row of A then det(A) = det(B)
1
Expert's answer
2020-09-23T17:35:51-0400

"\textsf{Let}\hspace{0.1cm}A =\begin{pmatrix} 
a_1 & a_2 & a_3\\
a_4 & a_5 & a_6\\
a_7 & a_8 & a_9
\end{pmatrix}\\

\mathrm{det} A = a_1(a_5(a_9) - a_6(a_8)) - a_2(a_4(a_9) - a_6(a_7)) + a_3(a_4(a_8) - a_5(a_7))\\

B = 
\begin{pmatrix} 
a_1 & a_2 & a_3\\
a_4 & a_5 & a_6\\
a_7 + 2a_1 & a_8 + 2a_2 & a_9 + 2a_3
\end{pmatrix}\\

\mathrm{det}B = a_1(a_5(a_9 + 2a_3) - a_6(a_8 + 2a_2)) - a_2(a_4(a_9 + 2a_3) - a_6(a_7 + 2a_1)) + a_3(a_4(a_8 + 2a_2) - a_5(a_7 + 2a_1)) =\\ a_1(a_5(a_9) - a_6(a_8)) - a_2(a_4(a_9) - a_6(a_7)) + a_3(a_4(a_8) - a_5(a_7)) + 2(a_1a_3a_5 - a_1a_2a_6 - a_2a_3a_4 + a_1a_2a_6 + a_2a_3a_4 - a_1a_3a_5) =\\ a_1(a_5(a_9) - a_6(a_8)) - a_2(a_4(a_9) - a_6(a_7)) + a_3(a_4(a_8) - a_5(a_7)) + 2(0) =\\ a_1(a_5(a_9) - a_6(a_8)) - a_2(a_4(a_9) - a_6(a_7)) + a_3(a_4(a_8) - a_5(a_7)) = \mathrm{det} A\\


\therefore \mathrm{det}B= \mathrm{det}A"